A cashier has four times as many nickels as quarter. She has $3.60. How many of each coin does she have?
Answer by ptaylor(2198)

Let x=number of quarters
Then 4x=number of nickels
Let's deal in pennies
Value of quarters=25x
Value of nickels=4*5x=20x
And we are told that the above amounts=360
So, our equation to solve is:
25x+20x=360 collect like terms
45x=360 divide each side by 45
x=8--------------------------------number of quarters
4x=4*8=32--------------------------------------number of nickels
CK
25*8+5*32=360
200+160=360
360=360
Does this help??-----ptaylor